OpenStudy (anonymous):

A red train traveling at 72 km/hr and a green train traveling at 144 km/hr are headed toward one another along a straight level track. When they are 950 meters apart, each enginneer sees the other's train and applies the brakes. The brakes decelerate each train at the rate of 1.0 m/s^2. Is there a colllision? If so, what is the speed of each train at impact? If not, what is the separtion between the trains when they stop??

OpenStudy (dls):

For red train: v^2-u^2=-2as 0-400=-2s s=200m Red train travels 200m.. For green train: v^2-u^2=-2as 0-1600=-2s s=800m Red train travels 800m.. Hence..they will collide because..800+200=1000m>950m(initial seperation)

OpenStudy (dls):

clear?

OpenStudy (anonymous):

but what was the speed of the trains when they collided?

OpenStudy (dls):

Red=>72km/hr=20m/s Green=>144km/hr=40m/s

OpenStudy (anonymous):

No i meant, when the two trains collided..

OpenStudy (dls):

0

OpenStudy (dls):

You can do this via Relative Velocity too

OpenStudy (dls):

That is a one line answer :)

OpenStudy (anonymous):

if the speed was 0, they would have stopped tho..

OpenStudy (dls):

We are only asked if they collide..so we assume final velocity=0

OpenStudy (anonymous):

haha it said "Is there a colllision? If so, what is the speed of each train at impact?"

OpenStudy (dls):

so 0 :)

OpenStudy (dls):

not sure..

OpenStudy (dls):

wait

OpenStudy (anonymous):

but 0 means they trains are stopped. so there won't be a collision and the answer would be "no".

OpenStudy (dls):

there you go o.O

OpenStudy (anonymous):

hmm.. what is 35.83m/s?

OpenStudy (dls):

final velocity of red train!

OpenStudy (anonymous):

so it actually becomes faster??

OpenStudy (anonymous):

even tho it's braking?

OpenStudy (dls):

u have the answer key?

OpenStudy (dls):

a=-1 not 1!

OpenStudy (dls):

v=u-at

OpenStudy (dls):

you have 3 things

OpenStudy (anonymous):

I know that they collide and red train stops

OpenStudy (anonymous):

I just need to know the speed of the green train at the point of the impact.

OpenStudy (dls):

Consider red train to be at rest and apply relative velocity u have 3things

OpenStudy (anonymous):

??

OpenStudy (dls):

what's the ans

OpenStudy (anonymous):

i don't know.......